OpenStudy (anonymous):

Simplify:

OpenStudy (anonymous):

Simplify:

OpenStudy (anonymous):

None of the above

ganeshie8 (ganeshie8):

\(\sqrt{10} \times \sqrt{8} = \sqrt{10\times 8} = \sqrt{2\times 5 \times 2^3} = \sqrt{16 \times 5} = 4\sqrt{5} \)